OpenStudy (anonymous):

Help!! I am doing Algebra I online this summer, and I don?t really get this question... can anyone explain? Which of the following points lie in the solution set to the following system of inequalities? y <_ x - 5 y <_-x - 4 (1, 10) (-1, 10) (10, 1) (1, -10) *MEDALS*

OpenStudy (across):

Try all choices by plugging in the given ordered pairs. You will find that: -10 <= (1) - 5 -10 <= - (1) - 4 (1, -10)

OpenStudy (anonymous):

(-1, -10) is correct. Graph both, and see which point is shaded twice.

OpenStudy (anonymous):

If a point is a solution to a set of equations or inequalities you just have to show that it satisfies them all. If you plug in the (x,y) pair you will either get a statement that is true: \[-10 \le -4\] Or a statement that is false: \[10 \le -4\] In order to satisfy the system it has to be true for all the equations/inequalities in the system.